by wieght, one fertilizer is 20% pottassium, 30% nitrogen, and 50% phosphorus. a second fertilizer has percents of 10, 30, and 60, respectively, and a third fertilizer has percents of 0, 20, 80, respectively.
How much of each must be mixed to get 200lb of fertilizer with percents of 12, 28.5, 59.5, respectively?

Answer by stanbon(75887)

by weight,
A fertilizer is 20% potassium, 30% nitrogen, and 50% phosphorus.
B fertilizer has percents of 10, 30, and 60,
C fertilizer has percents of 0, 20, 80, respectively.
------
How much of each must be mixed to get 200lb of fertilizer with percents of 12, 28.5, 59.5, respectively?
---
Equations:
Potassium Eq: 0.20a + 0.10b + 0*c = 0.12*200
Nitrogen Eq:: 0.30a + 0.30b + 0.20c = 0.285*200
Phosphorus Eq:0.50a + 0.60b + 0.80c = 0.595*200
-----------------------------------------------
I used a matrix method to get:
A = 70 lbs.
B = 100 lbs
C = 30 lbs.
==============
Cheers,
Stan H.